OpenStudy (anonymous):

Evaluate the expression. (6 – 2)[(2 + 3) + 4] A. 9 B. 13 C. 24 D. 36

OpenStudy (defiance):

Solve brackets first. Then Parenthesis inside brackets (If there is) 2+3 = 5 5 + 4 = 9 Solve the other parenthesis. 6 - 2 = 4 (4)9 \[4 \times 9 \] makes 36